OpenStudy (anonymous):

4x^2 + (y-2)^2 find dy/dx find the equation of the line tangent to the curve at point (0,6)

OpenStudy (anonymous):

dy/dx = 8x + 2(y-2)* dy/dx 2(y-2)* dy/dx - dy/dx = -8x dy/dx = -8x / [ 2(y-2) - 1] = -8x ( 2y -5)

OpenStudy (anonymous):

so that first part is equal to 16 so do i just do -8x(2y-5)=16 then solve for y?

OpenStudy (anonymous):

the 4x^2 + (y-2)^2 = 16 ?

OpenStudy (anonymous):

what was the original question?